# EvalNet: A Practical Toolchain for Generation and Analysis of Extreme-Scale Interconnects

### v3

@@ 3.7. Kautz Topologies
Both Undirected Kautz (KA) (Li et al., 2004) and Arrangement Network (Day and Tripathi, 1992) have strings assigned to each router. Then, links between routers are determined by similarities of the corresponding router strings. Kautz networks offer high resilience properties and offer an interesting alternative to other classes of interconnects.
3.8. Other Topologies
@@
EvalNet can be easily extended to model and evaluate interconnects from other domains, such as on-chip networks between different cores (Ahn et al., 2015; Besta et al., 2018), chiplets (Iff et al., 2023a, 2025, b), or FPGA interconnects (Cai et al., 2015), due to its modular and flexible architecture. The toolchain’s seamless integration with external simulators for packet-level and flow-level analysis ensures that new topologies, such as those designed for heterogeneous (Chang et al., 2012) or spatial computing environments (Gianinazzi et al., 2023; Baumann et al., 2024), can be evaluated with minimal effort, making EvalNet a versatile tool for broader interconnect research.
4. EVALUATING NETWORK TOPOLOGIES
@@
We next describe the considered measures used to evaluate a network topology. We analyzed a large body of related work to provide a comprehensive overview (Besta et al., 2021; Kim et al., 2008, 2007; Besta and Hoefler, 2014; Prisacari et al., 2013b; Yuan et al., 2014, 2013; Prisacari et al., 2014a, b; Kathareios et al., 2015; Prisacari et al., 2015; Chen et al., 2016; Prisacari et al., 2013a; Karacali et al., 2018; Sehery and Clancy, 2017; Kassing et al., 2017; Dally and Towles, 2003). We distinguish five classes of measures, used to analyze five aspects of a network: path diversity, bandwidth, latency, construction cost, and power consumption. All these metrics are also listed in Figure 3.
4.1. Path Diversity Measures
@@
Recent studies in high-performance interconnect design indicate that harnessing different forms of path diversity, especially non-minimal multipathing, is a key performance enabler, especially for state-of-the-art low-diameter networks such as SF or PF. This is because these networks – while ensuring low latency, high throughput, as well as low construction cost and power consumption – have only one shortest path between almost all router pairs (Besta et al., 2020, 2021). Hence, one cannot use traditional load-balancing mechanisms such as equal-cost multi-path routing (ECMP) (Hopps, 2000) for more performance, and must resort to multipath routing over non-minimal paths (Besta et al., 2020).
@@
To address this, EvalNet supports analyzing all recently proposed measures for path diversity. First, it enables computing counts of both minimal and non-minimal paths between any two routers. These two measures illustrate the potential for multipathing over both traditional and recent networks. Moreover, we use the path interference (PI) measure, which illustrates how multipathing changes when considering interactions between different pairs of routers (i.e., some parts of a given network may be more or less advantageous for multipathing than other parts of the same network). This enables deriving detailed insights into path diversity in the local network structure, facilitating the fine-grained design of routing protocols. Finally, EvalNet enables computing the total network load (TNL), which is an upper bound on the number of flows that a network can maintain without congestion; TNL constitutes the maximum supply of path diversity offered by a given topology. We define all these measures in more detail in Section 5 while analyzing their values for different topologies.
@@
Figure 4. (left) Histograms of shortest path lengths $l_{\text{min}}$. (right) Histograms of shortest path multiplicities $n_{\text{min}}$. Considered networks: DF, SF, and two HX variants with $D\in\{2,3\}$.
[…]
[Histograms of shortest path lengths and multiplicities]The figure shows in two subfigures the histograms of the shortest path lengths and multiplicities for four topologies: Dragonfly, Slim Fly and HyperX (diameter of 2 and 3). Four different server sizes are illustrated: 1.000, 10.000, 100.000 and 1.000.000. The shortest path lengths facilitate reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs. More router pairs in Slim Fly are connected with a 2-hop shortest path, than in HyperX (diameter of 2), for all topology sizes. This motivates more focus on making best-possible routing decisions over 2-hop paths in Slim Fly; 1-hop paths are negligible (especially for large networks). The shortest path multiplicities enable taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing. Nearly all router pairs in Slim Fly have only 1 shortest path between them, making non-minimal multipathing crucial. Most router pairs in HyperX have many shortest paths between them, and this ratio is growing with the number of servers, making ECMP a good option.
4.2. Throughput & Latency Measures
@@
Network latency is also obtained using packet-based simulations. The network latency is defined as the latency achieved for a very low injection rate that is close to zero. Thus, latency measures the end-to-end communication delay without congestion effects in the network. Moreover, the interfaced simulators enable delivering the average path length, i.e., the average number of router-router links traversed between any two communicated endpoints. This measure is useful because it is related to the network topological structure, i.e., one can see more clearly (than with latency) the impact of a routing design decision on the number of links that must be traversed.
4.3. Construction Cost & Power Consumption
@@
5. PATH DIVERSITY ANALYSIS
We present path diversity measures and conduct analyses in EvalNet, obtaining interesting insights about various interconnects. These measures are defined on pairs of routers (Besta et al., 2020). Extended results are provided in the Appendix. We start with basic definitions shared by all measures.
Definition 5.1.
@@ 5. PATH DIVERSITY ANALYSIS
Figure 5. The analysis of non-shortest path diversity and path interference.
[…]
[Histograms of multiplicities of non-shortest paths and path interferences for shortest & non-shortest paths as well as the path diversity between all router pairs and 2D histograms of path interferences for shortest & non-shortest paths]The figure analyses non-shortest path diversity and path interference in four subfigures. The compared topologies are Dragonfly, SlimFly and HyperX (diameter of 2 and 3). The multiplicities of non-shortest paths facilitate reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on routing design. HyperX with a diameter of 2 has very high path diversity that is also ’stable’ (20 for all router pairs) for all higher path lengths (l=3,4,5). This facilitates the design of complex adaptive routing that routes around hotspots, uses complex paths, et cetera. The analysis of path interferences for shortest & non-shortest paths facilitates reasoning about the performance of both minimal and non-minimal routing, and of scheduling of potential processes to different servers. Dragonfly, HyperX with a diameter of 3, and Slim Fly have almost no interference at two hops. Thus, one does not have to pay much attention to how to schedule pairs of communicating jobs, assuming they are executed on servers attached to routers separated by the distance l=2 (or l=1). The analysis of the path diversity between all router pairs facilitates reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on how to route. The majority of router pairs are connected by a number of different edge-disjoint paths that is larger than 1/3 of all their available network ports (indicated with circles and triangles). Such pairs do not require a lot of attention when selecting a path to route. The analysis of path interferences for shortest & non-shortest paths acilitates reasoning about the performance of both minimal and non minimal routing, and of scheduling of potential processes to different servers is presented in the last subfigure as 2D histograms.
5.1. Diversity of Shortest Paths
First, we consider the lengths and multiplicities of shortest paths. Understanding the former facilitates reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs. Knowing the latter, on the other hand, enables taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing (Besta et al., 2020, 2021).
Definition 5.4.
@@ 5.1. Diversity of Shortest Paths
Definition 5.5.
The shortest path multiplicity between $s,t\in V$ counts the number of shortest paths between $s$ and $t$ and is defined as
@@
We provide example results in Figure 4, for both lengths of shortest paths (the left panel) and for their multiplicities (the right panel), for different sizes of four selected networks (indicated with different colors). The left panel illustrates the rich information about lengths of shortest paths. For example, most pairs are connected with a shortest path having two hops. Specifically, more router pairs in SF are connected with a 2-hop shortest path, than in HX2, for all topology sizes. This motivates more focus on making best-possible routing decisions over 2-hop paths in SF; 1-hop paths are negligible (especially for large networks). The right plot shows that, for most router pairs, the multiplicity of DF, HX2, and SF is low ($n_{\text{min}}\leq 2$) compared to HX3. We also observe that nearly all router pairs in SF have only 1 shortest path between them, making non-minimal multipathing crucial. Still, note that DF and SF host some outlier router pairs connected with large counts of shortest paths. Finally, most router pairs in HX have many shortest paths between them, and this ratio is growing with $N$, making ECMP a good option.
5.2. Diversity of Non-Shortest Paths
@@
We show example results in Figure 5. The top-left panel contains the histograms of $c_{l}$ at various lengths for DF, SF, and two HX instances. The values of $c_{l}$ of DF are in general lower than in the low dimensional HX instances (HX2 and HX3) and SF. Accordingly, $c_{l}$ of the HX and SF instances attains the network radix at lower lengths. HX2 has very high path diversity that is also “stable” (20 for all router pairs) for all higher path lengths ($l=3,4,5$). This facilitates the design for complex adaptive routing that routes around hotspots, uses complex paths, et cetera. Next, DF has relatively large diversity of paths for $l=5$, suggesting similar insights as for HX2, but the count is 11, so less than 20 by nearly 50%. Then, for $l=4$ or $l=3$, there is quite some diversity but the counts for different pairs may differ significantly, requiring attention when developing routing.
@@ 5.3. Path Interference
5.4. Computing Path Diversity Efficiently
@@
In EvalNet, we primarily use algebraic algorithms to compute path diversity metrics, as they offer better scalability and parallelizability compared to purely combinatorial approaches, particularly for large, low-diameter networks (Besta et al., 2017; Solomonik et al., 2017; Solomonik and Demmel, 2011; Solomonik et al., 2014). For example, for counting the number of paths of fixed length between all node pairs, we use adjacency matrix exponentiation, where the $\ell$-th power of the adjacency matrix $A$ gives the number of $-\ell$ paths via $(A^{\ell}){i,j}$. This computation is efficient for small diameters and can be adapted for next-hop tracking using semiring-based variants. For length-restricted connectivity, we extend the algebraic path propagation algorithm of Cheung et al. (Cheung et al., 2011). For other metrics, we also explore max-flow-based approaches such as Push-Relabel and Ford-Fulkerson with BFS path constraints (used for sampled length-limited disjoint path analysis), as well as classic Gomory-Hu trees for all-pairs edge connectivity. While these combinatorial algorithms serve as correctness baselines or targeted tools for small-scale samples, EvalNet’s core design centers on the algebraic methods due to their efficiency, extensibility, and compatibility with parallel computation.
@@
Figure 6. (left) Heatmap showing the Maximum Achievable Throughout (MAT), obtained with TopoBench, for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern. (right) Packet-based simulation runtimes [cycles] with Booksim2, for networks of 50k endpoints, for two injection rates, 0.01 and 0.5. As expected, networks with lower diameter usually provide lower latency (load=0.01). For load=0.5, the network global bandwidth starts to play an important role; here, DF has higher latency than FT3 because it has much lower global bandwidth.
[…]
[Figure is fully described in the text]Figure is fully described in the text.
6. FLOW-LEVEL SIMULATIONS
@@
Flow-level simulators enable evaluation of large-scale systems, at a cost of a coarse design that may hinder realistic insights into performance. Flow-level simulations assume a continuous, real-valued flow model. Each flow has a rate $\mu\in\mathbb{R^{+}}$. The sum of all flows on each link must not exceed the link capacity. Without any additional constraints, this leads to constrained max-flow solutions in the topology graph. Other constrains (e.g., on the routes that can be taken by specific flows) lead to more realistic upper-bounds for performance of various routing schemes on specific topologies. However, since these models do not represent time accurately, they are unable to estimate latency or provide realistic insights into the impact of routing choices on flow control. As such, they are a useful tool for initial insights when designing a network.
@@
For flow simulations, we provide an interface to TopoBench, an evaluation tool (Jyothi et al., 2016) that uses linear programming (LP) to derive MAT. Our TopoBench version supports state-of-the-art layered routing schemes such as FatPaths (Besta et al., 2020) or extended layered routing (Blach et al., 2024), built specifically for enabling non-minimal multi-path routing. Example results are in the left part of Figure 6, where we analyze a variant of FatPaths that offers non-minimal multipath routing. That part contains a heatmap showing MAT for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern. “#R” (Y axis) as well as “#P” (X axis) are parameters of the FatPaths routing scheme; they can be tuned for a specific topology. Usually, the higher #R is, the better the throughput becomes (because more layers entail more potential for multipathing (Besta et al., 2020)). Simultaneously, higher #R indicates more HW costs because one needs larger forwarding tables. Then, higher #P does not always enhance MAT (as confirmed by the results), because certain value ranges may cause larger path interference across routing layers, ultimately diminishing performance (Besta et al., 2020). Overall, for certain configurations, FatPaths on SF achieves very high MAT.
7. PACKET-LEVEL SIMULATIONS
@@
7.1. Which Simulator to Use?
A large number of simulators exist, and it is challenging to select the one best suited for one’s purposes. We analyze a broad selection of available simulators, and show that none scale to the desired sizes of hundreds of thousands of servers. The overview is in Table 2. We analyzed their codebases, investigating which (1) facilitate modifications that could enhance scalability, (2) are easy to use and extend without requiring distributed-memory infrastructure support, and (3) cover a large number of protocols and settings. Based on that analysis, we selected two simulators: htsim (Raiciu et al., 2010) for lossy TCP/IP stacks and Booksim2 (Jiang et al., 2013) for lossless HPC architectures. Other simulators do not meet these criteria. For example, SST and ROSS natively target HPC clusters and we concluded that they are relatively complex to set up, use, and extend on low-cost PCs.
htsim The htsim simulator (Raiciu et al., 2010) provides a lightweight infrastructure for lossy networks. It only models the transport layer: there is no model for links or routers. Instead, the route for each packet is pre-computed as a sequence of queues that the packet will pass through. Such a route specification is attached to the packet. Each queue has a finite service rate, which models the link capacity. Due to this design, htsim is highly flexible, and new topologies can be added straightforwardly: the topology only affects the route computation, which is explicitly called for every flow during its setup. The disadvantage is that many adaptive routing schemes cannot be modeled, since there is no per-router state that could affect the routing. Furthermore, all flows are initialized before the first event is processed, and all routes are kept in memory. EvalNet offers a seamless interface to htsim. We select htsim as it targets TCP stacks; it supports by default modern TCP-based schemes such as NDP (Handley et al., 2017), and it can be easily extended with new routing schemes.

### v4

@@ III-G Kautz Topologies
Both Undirected Kautz (KA) [70] and Arrangement Network [30] have strings assigned to each router. Then, links between routers are determined by similarities of the corresponding router strings. Kautz networks offer high resilience properties and are an interesting alternative to other classes of interconnects.
III-H Other Topologies
@@
EvalNet can be easily extended to model and evaluate interconnects from other domains, such as on-chip networks between different cores [2, 14], chiplets [50, 52, 51], or FPGA interconnects [20], due to its modular and flexible architecture. The toolchain’s seamless integration with external simulators for packet-level and flow-level analysis ensures that new topologies, such as those designed for heterogeneous [24] or spatial computing environments [40, 10], can be evaluated with minimal effort, making EvalNet a versatile tool for broader interconnect research.
IV EVALUATING NETWORK TOPOLOGIES
@@
We next describe the considered measures used to evaluate a network topology. We analyzed a large body of related work to provide a comprehensive overview [13, 63, 64, 15, 89, 107, 108, 85, 86, 60, 87, 25, 88, 58, 93, 59, 29]. We distinguish five classes of measures, used to analyze five aspects of a network: path diversity, bandwidth, latency, construction cost, and power consumption. All these metrics are also listed in Figure 3.
IV-A Path Diversity Measures
@@
Recent studies in high-performance interconnect design indicate that harnessing different forms of path diversity, especially non-minimal multipathing, is a key performance enabler, especially for state-of-the-art low-diameter networks such as SF or PF. This is because these networks – while ensuring low latency, high throughput, as well as low construction cost and power consumption – have only one shortest path between almost all router pairs [17, 13]. Hence, one cannot use traditional load-balancing mechanisms such as equal-cost multi-path routing (ECMP) [47] for more performance, and must resort to multipath routing over non-minimal paths [17].
@@
To address this, EvalNet supports analyzing all recently proposed measures for path diversity. First, it enables computing counts of both minimal and non-minimal paths between any two routers. These two measures illustrate the potential for multipathing over both traditional and recent networks. Moreover, we use the path interference (PI) measure, which illustrates how multipathing changes when considering interactions between different pairs of routers (i.e., some parts of a given network may be more or less advantageous for multipathing than other parts of the same network). This enables deriving detailed insights into path diversity in the local network structure, facilitating the fine-grained design of routing protocols. Finally, EvalNet enables computing the total network load (TNL), which is an upper bound on the number of flows that a network can maintain without congestion; TNL constitutes the maximum supply of path diversity offered by a given topology. We define all these measures in more detail in Section V while analyzing their values for different topologies.
@@
Figure 4: (left) Histograms of shortest path lengths $l_{\text{min}}$. (right) Histograms of shortest path multiplicities $n_{\text{min}}$. Considered networks: DF, SF, and two HX variants with $D\in\{2,3\}$.
IV-B Throughput & Latency Measures
@@
Network latency is also obtained using packet-based simulations. The network latency is defined as the latency achieved for a very low injection rate that is close to zero. Thus, latency measures the end-to-end communication delay without congestion effects in the network. Moreover, the interfaced simulators enable delivering the average path length, i.e., the average number of inter-router links traversed between any two communicated endpoints. This measure is useful because it is related to the network topological structure, i.e., one can see more clearly (than with latency) the impact of a routing design decision on the number of links that must be traversed.
IV-C Construction Cost & Power Consumption
@@
V PATH DIVERSITY ANALYSIS
We present path diversity measures and conduct analyses in EvalNet, obtaining interesting insights about various interconnects. These measures are defined on pairs of routers [17]. Extended results are provided in the Appendix. We start with basic definitions shared by all measures.
Definition V.1.
@@ V PATH DIVERSITY ANALYSIS
Figure 5: The analysis of non-shortest path diversity and path interference.
V-A Diversity of Shortest Paths
First, we consider the lengths and multiplicities of shortest paths. Understanding the former facilitates reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs. Knowing the latter, on the other hand, enables taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing [17, 13].
Definition V.4.
@@ V-A Diversity of Shortest Paths
Definition V.5.
The shortest path multiplicity between $s,t\in V$ is the number of shortest paths between $s$ and $t$ and is defined as
@@
We provide example results in Figure 4, for both lengths of shortest paths (the left panel) and for their multiplicities (the right panel), for different sizes of four selected networks (indicated with different colors). The left panel illustrates the rich information about lengths of shortest paths. For example, most pairs are connected with a shortest path having two hops. Specifically, more router pairs in SF are connected with a 2-hop shortest path than in HX2 for all topology sizes. This motivates a focus on making best-possible routing decisions over 2-hop paths in SF; 1-hop paths are negligible (especially for large networks). The right panel shows that, for most router pairs, the multiplicity of DF, HX2, and SF is low ($n_{\text{min}}\leq 2$) compared to HX3. We also observe that nearly all router pairs in SF have only 1 shortest path between them, making non-minimal multipathing crucial. Still, note that DF and SF contain some outlier router pairs connected with large counts of shortest paths. Finally, most router pairs in HX have many shortest paths between them, and this ratio is growing with $N$, making ECMP effective.
V-B Diversity of Non-Shortest Paths
@@
We show example results in Figure 5. The top-left panel contains the histograms of $c_{l}$ at various lengths for DF, SF, and two HX instances. The values of $c_{l}$ of DF are in general lower than in the low dimensional HX instances (HX2 and HX3) and SF. Accordingly, $c_{l}$ of the HX and SF instances attains the network radix at lower lengths. HX2 has very high path diversity that is also “stable” (20 for all router pairs) for all higher path lengths ($l=3,4,5$). This facilitates the design for complex adaptive routing that routes around hotspots and uses complex paths. Next, DF has relatively large diversity of paths for $l=5$, suggesting similar insights as for HX2, but the count is 11, so less than 20 by nearly 50%. Then, for $l=4$ or $l=3$, there is quite some diversity but the counts for different pairs may differ significantly, requiring attention when developing routing.
@@ V-C Path Interference
V-D Computing Path Diversity Efficiently
@@
In EvalNet, we primarily use algebraic algorithms to compute path diversity metrics, as they offer better scalability and parallelizability compared to purely combinatorial approaches, particularly for large, low-diameter networks [16, 95, 97, 96]. For example, for counting the number of paths of fixed length between all node pairs, we use adjacency matrix exponentiation, where the $\ell$-th power of the adjacency matrix $A$ gives the number of length-$\ell$ paths via $(A^{\ell}){i,j}$. This computation is efficient for small diameters and can be adapted for next-hop tracking using semiring-based variants. For length-restricted connectivity, we extend the algebraic path propagation algorithm of Cheung et al. [26]. For other metrics, we also explore max-flow-based approaches such as Push-Relabel and Ford-Fulkerson with BFS path constraints (used for sampled length-limited disjoint path analysis), as well as classic Gomory-Hu trees for all-pairs edge connectivity. While these combinatorial algorithms serve as correctness baselines or targeted tools for small-scale samples, EvalNet’s core design centers on the algebraic methods due to their efficiency, extensibility, and compatibility with parallel computation.
@@
Figure 6: (left) Heatmap showing the Maximum Achievable Throughput (MAT), obtained with TopoBench, for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern. (right) Packet-based simulation runtimes [cycles] with Booksim2, for networks of 50k endpoints, for two injection rates, 0.01 and 0.5. As expected, networks with lower diameter usually provide lower latency (load=0.01). For load=0.5, the network global bandwidth starts to play an important role; here, DF has higher latency than FT3 because it has much lower global bandwidth.
VI FLOW-LEVEL SIMULATIONS
@@
Flow-level simulators enable evaluation of large-scale systems, at a cost of a coarse design that may hinder realistic insights into performance. Flow-level simulations assume a continuous, real-valued flow model. Each flow has a rate $\mu\in\mathbb{R^{+}}$. The sum of all flows on each link must not exceed the link capacity. Without any additional constraints, this leads to constrained max-flow solutions in the topology graph. Other constraints (e.g., on the routes that can be taken by specific flows) lead to more realistic upper-bounds for performance of various routing schemes on specific topologies. However, since these models do not represent time accurately, they are unable to estimate latency or provide realistic insights into the impact of routing choices on flow control. As such, they are a useful tool for initial insights when designing a network.
@@
For flow simulations, we provide an interface to TopoBench, an evaluation tool [57] that uses linear programming (LP) to derive MAT. Our TopoBench version supports state-of-the-art layered routing schemes such as FatPaths [17] or extended layered routing [18], built specifically for enabling non-minimal multi-path routing. Example results are in the left part of Figure 6, where we analyze a variant of FatPaths that offers non-minimal multipath routing. That part contains a heatmap showing MAT for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern. “#R” (Y axis) as well as “#P” (X axis) are parameters of the FatPaths routing scheme; they can be tuned for a specific topology. Usually, the higher #R is, the better the throughput becomes (because more layers entail more potential for multipathing [17]). Simultaneously, higher #R indicates more HW costs because one needs larger forwarding tables. Then, higher #P does not always enhance MAT (as confirmed by the results), because certain value ranges may cause larger path interference across routing layers, ultimately diminishing performance [17]. Overall, for certain configurations, FatPaths on SF achieves very high MAT.
VII PACKET-LEVEL SIMULATIONS
@@
VII-A Which Simulator to Use?
A large number of simulators exist, and it is challenging to select the one best suited for one’s purposes. We analyze a broad selection of available simulators, and show that none scale to the desired sizes of hundreds of thousands of servers. The overview is in Table II. We analyzed their codebases, investigating which (1) facilitate modifications that could enhance scalability, (2) are easy to use and extend without requiring distributed-memory infrastructure support, and (3) cover a large number of protocols and settings. Based on that analysis, we selected two simulators: htsim [91] for lossy TCP/IP stacks and Booksim2 [56] for lossless HPC architectures. Other simulators do not meet these criteria. For example, SST and ROSS natively target HPC clusters and we concluded that they are relatively complex to set up, use, and extend on low-cost PCs.
htsim The htsim simulator [91] provides a lightweight infrastructure for lossy networks. It only models the transport layer: there is no model for links or routers. Instead, the route for each packet is pre-computed as a sequence of queues that the packet will pass through. Such a route specification is attached to the packet. Each queue has a finite service rate, which models the link capacity. Due to this design, htsim is highly flexible, and new topologies can be added straightforwardly: the topology only affects the route computation, which is explicitly called for every flow during its setup. The disadvantage is that many adaptive routing schemes cannot be modeled, since there is no per-router state that could affect the routing. Furthermore, all flows are initialized before the first event is processed, and all routes are kept in memory. EvalNet offers a seamless interface to htsim. We select htsim as it targets TCP stacks; it supports by default modern TCP-based schemes such as NDP [42], and it can be easily extended with new routing schemes.